OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (anonymous):

Well, first, solve your 2x - 3y = 12 equation for y, i.e., make it in conventional slope-intercept form. Then, once you find your slope (m), take the negative reciprocal (-1/m) and that will be the slope of the line perpendicular. If it passes through point (2,6), so you can plug those values into your new equation and solve for B (the y-intercept), and then you will be able to fully-build your new equation. Try this and tell me what you get!

OpenStudy (***[isuru]***):

first transform the given equation to the general format of y = mx + c where m = slope c = intercept and u have to remember this also when there r 2 lines which r perpendicular to each other , if the slope of one line is "m" then the slope of the other is " (-1/m)" so... moving forward.. when u transform 2x - 3y = 12 into y = mx + c format u will get... 3y = 2x -12 y = (2/3)x - 4 now remember what i mention about the slope.. if the slope of one is "m" then the slope of the other one is " -1/m" according to that the slope of the line we want to draw will be (-3/2) then we can write the equation of the line we want as.. y= (-3/2) x + c let's consider the given points (2 , 6) by substituting them in the equation we get 6 = (-3/2)2 + c then c will be c = 9 therefor the equation of the line will be y = (-3/2)x + 9

OpenStudy (anonymous):

thanks both of you for your explinations on how to solve them :)